OpenStudy (anonymous):

The sum of two numbers is  26.  The first number is  3/10  of the second number. What are the numbers?

OpenStudy (anonymous):

Hi welcome to Openstudy! let x be the first number and y be the 2nd number x + y = 26 x = 3/10y since it says "The first number is  3/10  of the second number" solve for the two unknown numbers by combining the two equations we have: x+ y = 26 3/10 y + y =26 can you simplify this to solve for y? i'll check it don't worry, just try :)

OpenStudy (anonymous):

20?

Nnesha (nnesha):

:-)yep

Nnesha (nnesha):

ahem why did you delete that ?? O.o

OpenStudy (anonymous):

yes great :) so now let's solve for x, using x= 3/10 y plug in the y value that you calculated in x = 3/10 y and tell us what you will get

OpenStudy (anonymous):

delete what? ? O-o?

Nnesha (nnesha):

*secret*

OpenStudy (anonymous):

lol okay :3

OpenStudy (anonymous):

@idiotXatXmath what are we going to get if replace y with 20 in x= 3/10 y? it will look like this: x = 3/10 (20) which is equal to?

OpenStudy (anonymous):

my bad 6y

OpenStudy (anonymous):

yes! x will be six therefore the unknown numbers that you are looking for are 20 and 6 ^_^ great job @idiotXatXmath !

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anytime :P